# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 4.2.2 Peer LLMs as Verifiers
A growing line of work leverages LLMs as external evaluators, distinct from the generator itself. For example, FlexiVe (zhong2025solvedetectverifyinferencetimescalingflexible) implements a solve–detect–verify pipeline, while process-reward frameworks (choudhury2025processrewardmodelsllm) and diverse verifier tree search use verifier models to score intermediate hypotheses and control expansions. Stability can be improved by ensemble verification (wang2025dontlosttreesstreamlining).
The CRPO (2025crpo) exemplifies this approach by using an LLM’s general linguistic knowledge as a feedback signal for guiding logical consistency and language fluency. Similarly, the LLM-as-a-Judge (zheng2023judgingllmasajudgemtbenchchatbot) paradigm formalizes this idea: one model (or ensemble (chen2025multiagentasjudgealigningllmagentbasedautomated)) evaluates hypotheses produced by another, effectively operationalizing externalized language-model judgment as search heuristics.
More advanced methods first train specialized evaluators to capture problem-specific criteria before deploying them as verifiers. For example, DeepResearcher in the Cycle-Researcher system (weng2025cycleresearcher) relies on a trained verifier model (DeepReviewer) to judge intermediate steps and direct search trajectories. Other frameworks like (li2024learninggenerateresearchidea) similarly builds a reward dataset and trains three reward models to guide idea generation dynamically, allowing fine-grained control over different evaluation dimensions. These approaches are often positioned as scalable, automated substitutes for human feedback, especially in domains where expert evaluation is costly. Peer evaluators are more flexible than self-confidence, since they may be domain-specialized and provide rationales.
However, they remain vulnerable to reward hacking: generators can learn to exploit systematic weaknesses in verifier heuristics, producing outputs that appear highly rated without being genuinely correct or useful. Recent studies in preference optimization and RLHF have shown that models can overfit to their evaluators, generating verbose, sycophantic, or superficially persuasive responses that maximize reward signals rather than substantive quality (laidlaw2025correlatedproxiesnewdefinition; miao2024inform). These findings highlight that while peer evaluators extend the evaluative horizon, their reliability depends on careful design to avoid exploitation and collapse of evaluation diversity. While effective as evaluators and filters, feedback sources differ fundamentally from collaborative multi-agent setups. As outlined in Section 5, multi-agent frameworks engage additional LLMs as co-creative partners, enabling them not only to assess but also to propose, contest, and refine ideas in a shared search process.
@@ 4.2.3 Simulators and Real Environments
Simulators serve as semi-grounded environments that test the outputs of LLMs in dynamic, feedback-rich contexts. These can be physical simulations, digital twins, or even agent-based societies. Simulators and real environments provide feedback through interactive trial-and-error, aligning closely with notion of online search(russell2016aima). Unlike offline planning, where a complete solution is computed in advance, online search emphasizes interleaving generation with execution: the agent proposes an action, observes the environment’s response, and adapts its trajectory accordingly. This makes simulators and lab-in-the-loop systems natural analogues, as they allow LLM-driven agents to iteratively refine hypotheses or strategies based on grounded feedback from physical laws, empirical trials, or synthetic environments.
Robin (ghareeb2025robinmultiagentautomatingscientific) uses a “lab-in-the-loop” design, where hypotheses are tested via in silico simulators (or even robotic wet-labs), and the resulting outputs are looped back into the ideation process as reward signals. For example, a gene-editing strategy might be generated, simulated for efficacy, and then refined based on empirical response.
Similarly, AtomAgents (ghafarollahi2024atomagentsalloydesigndiscovery) combines LLMs with physics-based alloy simulators: agents propose material compositions, query simulators for predicted properties like tensile strength, and prioritize candidates for further iteration based on performance metrics.
In more abstract contexts, CASEVO (jiang2024casevocognitiveagentssocial) simulates agent societies debating political or scientific stances, tracking how persuasive interactions shift beliefs—modeling ideation in socially complex domains.
Recent robotics work demonstrates the coupling of simulators and real execution. SayCan (ahn2022icanisay) grounds LLM-generated high-level commands in an affordance model and low-level robot controller, combining planning tools with real robot feedback. RoboTool (xu2023creativerobottooluse) extends this to creative tool use, where LLMs propose robot code tested in both simulators and physical hardware. Voyager (wang2023voyageropenendedembodiedagent) highlights how open-ended exploration in Minecraft enables the synthesis of reusable tools, iteratively evaluated in a simulator to accumulate a growing library of skills.
On the evaluation side, torne2024reconcilingrealitysimulationrealtosimtoreal analyze the fidelity of simulator-to-reality transfer, explicitly comparing robot manipulation policies in simulation versus physical execution. Similarly, Large-Scale Simulation of LLM-Driven Generative Agents stress-tests agentic LLMs in synthetic societies, showing how simulation feedback can approximate large-scale environment signals for long-running scientific ideation (piao2025agentsocietylargescalesimulationllmdriven). As shown in the Scenario 4.2.6, the simulator evaluates stability and feasibility of candidate designs.
@@ 4.2.4 External Tools and APIs
External tools expand the feedback horizon by integrating APIs, online search, and domain-specific databases. The Toolformer architecture (schick2023toolformer) equips LLMs with the capability to query external resources dynamically, enabling verification and augmentation of generated content. Chemical synthesis platforms like ChemCrow (bran2023chemcrow) leverage cheminformatics tools and external data sources to validate reaction pathways proposed by the model. Similarly, HuggingGPT (shen2023hugginggptsolvingaitasks) coordinates a wide range of specialized AI models hosted on the Hugging Face platform, allowing an LLM to act as a controller that routes tasks such as vision, speech, and text analysis to appropriate external tools, demonstrating how tool integration can expand an LLM’s capabilities beyond text-only reasoning.
Recent evaluations focus directly on LLM–tool interaction. MINT (wang2024mintevaluatingllmsmultiturn) systematically studies multi-turn tool usage, showing how feedback from tool calls and natural language guidance improves decision-making. Voyager’s open-ended tool synthesis also bridges this category, where tools (skills, functions) are generated internally but then tested through simulators and APIs.
By incorporating external tools, LLMs move beyond isolated language generation toward grounded, multi-modal reasoning informed by up-to-date, external knowledge. This shift transforms them from pure text predictors into decision-theoretic agents augmented with specialized oracles. This feedback source and its alternatives are examined in Scenario 4.2.6.
@@ 4.2.5 Scientific Priors and Domain Rules
SCaSML (patel2025scasml) integrates physical laws such as energy conservation to constrain generation, ensuring outputs remain within scientifically plausible bounds without requiring explicit simulation feedback. Similarly, methods like R³V (cheng2024r3v) embed domain logic checkpoints to detect and correct illogical reasoning mid-generation. The Scenario 4.2.6 illustrates how scientific priors enforce fundamental constraints and ensure theoretical validity. Domain rules are most effective in well-formalized fields (e.g., physics, chemistry, mathematics) where symbolic constraints or laws are clear and non-negotiable. Their strength is reliability: they are resistant to reward hacking and scale well. However, their rigidity limits flexibility in ill-defined or emergent domains, where strict rules may exclude innovative or unconventional ideas.
@@ 4.2.6 Human-in-the-Loop Evaluation
Platforms like (he2024collaborativeintelligencesequentialexperiments) integrate human expertise directly into the scientific discovery process by allowing researchers to iteratively guide and critique model-generated hypotheses. IdeaSynth (pu2025ideasynth) and Scideator (radensky2024scideator) provide interactive environments for hypothesis refinement, allowing researchers to blend and adapt elements from prior work. Chain-of-Table (lin2024chain) offers a mixed-initiative framework where human evaluation is combined with structured scientific tables and knowledge bases, balancing creativity with formal rigor. IRIS (feng-etal-2025-iris), though originally designed for self-reflection, also incorporates human evaluators in tasks requiring subjective judgment of novelty and coherence. These human-in-the-loop interactions enrich the feedback loop, ensuring that the discovery process not only leverages computational power for exploration but also incorporates experiential insights that are difficult to encode explicitly, ultimately leading to more reliable and interpretable scientific advancements
@@
Scenario 5: Contrasting Evaluation feedback sources To illustrate how different feedback sources shape ideation, consider an LLM that proposes a novel photo-catalyst for converting CO2 into useful chemicals under sunlight. The idea is passed through several evaluators, each providing a distinct form of guidance: Simulator: The design is tested in a simulation engine, which shows promising conversion rates but also reveals that the material becomes unstable at high temperatures. The simulator narrows the search space to only variants that remain physically viable. External database: The LLM queries scientific databases and patents. It finds no direct precedent for the exact design (a sign of novelty), but flags that similar catalysts once produced toxic byproducts. This feedback shifts exploration toward safer material choices. Scientific priors: Symbolic scientific rules check the proposed reaction. They detect a violation of basic energy conservation, forcing the model to reformulate the mechanism in a physically consistent way. Priors act as hard constraints that keep the search scientifically valid. Human expert: A human expert reviews the surviving ideas. Beyond technical feasibility, she considers manufacturability, cost, and environmental impact, and directs the system toward designs using abundant, scalable materials. Together, these perspectives highlight how different feedback sources contribute complementary strengths: simulators ground ideas in empirical behavior, external tools provide contextual knowledge, priors ensure theoretical soundness, and humans add judgment and broader vision.
@@ 4.3 Levels of Abstraction
Inference-time scaling methods differ not only in their strategies but also in the level of abstraction at which search is conducted. From a theoretical perspective, abstraction provides a means of reducing or restructuring the problem space, thereby enabling broader exploration but complicating verification (zucker2003abstraction; nakar2023abstraction). Recent LLM-based methods explicitly operationalize this trade-off. For instance, wang2023hypothesis show that separating hypothesis generation at an abstract level from concrete program instantiation improves performance on reasoning tasks, while singhal2024conceptsearch demonstrate that concept-level scoring functions allow more efficient traversal of search spaces than surface-level metrics. Together, these studies suggest that higher levels of abstraction broaden the search space and bring the process closer to exploratory creativity, but at the cost of increased difficulty in assessing feasibility and correctness.
@@ 4.3.1 Hypothesis-Level Ideation
For instance, MAGIC (wang2025magic) generates speculative hypotheses in genomics and systems biology, ranking them by plausibility and novelty heuristics. VirSci (johnson2024virsci) applies multi-agent debate to refine hypotheses through peer critique. IRIS (feng-etal-2025-iris) adds internal reflection, restructuring ideas mid-generation to avoid incoherence or excessive derivativeness. These systems act as cognitive companions: their outputs are inspirations requiring downstream verification by humans or automated evaluators. An example of hypothesis-level ideation is illustrated in Scenario 4.3.1.
Scenario 6: Hypothesis-Level Ideation (Mathematical Reasoning) An LLM proposes the conjecture: “For any odd prime $p$, the sum of quadratic residues modulo $p$ equals $\tfrac{p(p-1)}{4}$.” This is a purely abstract claim without proof. The model explores nearby conjectures (e.g., cubic residues, alternate modulus conditions). Evaluation requires external proof attempts or human mathematicians. Search style: Local search over conceptual variants of conjectures.
@@ 4.3.2 Program-Level Ideation
Program-Aided Language models (PAL) (gao2023pal) exemplify this principle by translating natural-language reasoning into Python programs executed externally, outperforming direct chain-of-thought reasoning.HuggingGPT (shen2023hugginggptsolvingaitasks), Toolformer (schick2023toolformer), and Explanations as Programs (zhang2023explanations**) similarly teach models to output modular, debuggable tool pipelines. In science, LLM-SR (shojaee2024llmsr) refines symbolic equations by querying solvers for constraint violations, ChemCrow (bran2023chemcrow) plans synthesis routes via cheminformatics APIs, and CodeScientist (codescientist2024**) generates executable lab protocols for simulators or robotic wet-labs.
Ideation at the program level enables the use of simulators and other external tools by integrating execution directly into the generative loop. An example of program-level ideation is illustrated in Scenario 4.3.2. This integration offers key benefits: programs are structured and interpretable, their execution provides robustness against unsupported claims, and hallucination is reduced as hypotheses must be empirically validated through execution.
Scenario 7: Program-Level Ideation (Mathematical Reasoning) Instead of stating a conjecture, the LLM outputs a structured proof plan: 1. Represent quadratic residues as $a^{2}\bmod p$. 2. Pair each $a^{2}$ with $(p-a)^{2}$. 3. Show these pairs sum to pp. 4. Derive a closed-form expression for the total sum. Each step can be checked in a proof assistant or tested computationally. Search style: Tree or beam search, where alternative proof strategies (induction vs. direct algebraic argument) form different branches.
@@ 4.3.3 Meta-Level Search and Self-Adaptation
A third, emerging frontier is meta-level search, where the object of optimization is not the hypothesis or program but the agent’s own configuration and interaction with its environment. Here, systems explore changes to their prompting strategies, exploration parameters, reward weightings, or even architectural components, adapting the search process itself. An example of meta-level ideation is illustrated in Scenario 4.3.3.
The Darwin Gödel Machine (zhang2025darwin) exemplifies self-modification: instead of emitting domain-level programs, it evolves its own reasoning code to improve performance. More broadly, meta-search also includes adaptive inference strategies—e.g., adjusting beam widths, modifying exploration–exploitation trade-offs, or re-balancing between simulator calls and heuristic priors. Such mechanisms parallel meta-reasoning and utility-based agent adaptation, where the agent learns not only how to act, but how to search. Another notable example is the Meta-review agent in Google’s AI co-scientist system (gottweis2025aicoscientist). This agent synthesizes feedback from various agents, identifies recurring issues, and generates comprehensive critiques and research overviews. The Meta-review agent compiles feedback across iterations to adjust the agents’ prompts, ensuring continual improvement.
@@
Scenario 8: Meta-Level Ideation (Mathematical Reasoning) The LLM reconfigures its own reasoning process rather than directly producing conjectures or proofs: – Switch from a “direct proof-first” mode to “counterexample-search-first” mode. – Invoke external theorem provers to prune invalid proof paths. – Deploy multiple agents specialized in algebraic manipulation vs. computational experiments. Here, the search space is over reasoning strategies themselves. Search style: Evolutionary or reinforcement-style optimization over proof-generation policies.
@@
Table 1: Comparison of scientific inference time scaling methods, grouped by search mechanism, feedback source and abstraction level
@@
Method Search Mechanism Feedback Source Abstraction Level Explanation
@@
Self-Refine(liu2023selfrefine) Sequential Refinement Internal confidence Hypothesis-Level Employs self-critiques iteratively to refine output.
@@
PANEL(li2025PANEL) Sequential Refinement Internal confidence Hypothesis-Level Refines ideas stepwise by generating critiques at each reasoning step.
@@
ResearchAgent(baek2024researchagent) Sequential Refinement Peer agents Hypothesis-Level Simulates peer-review with multiple agents providing feedback.
@@
CriticAL(li2024critical) Sequential Refinement Internal confidence Hypothesis-Level Uses statistical testing to criticize and improve reasoning.
@@
SCaSML(patel2025scasml) Sequential Refinement Scientific domain rules Program-Level Injects physical laws into inference and refines via Monte Carlo defect correction.
@@
Robin(ghareeb2025robinmultiagentautomatingscientific) Sequential Refinement Simulator Hypothesis-Level Refines ideas based on in-lab or in-silico experiment feedback.
AtomAgents(ghafarollahi2024atomagentsalloydesigndiscovery) Sequential Refinement Simulator Hypothesis-Level Refines alloy compositions through simulator feedback.
@@
IdeaSynth(pu2025ideasynth) Sequential Refinement Human + Tool Hypothesis-Level Researchers iteratively refine idea facets with human guidance.
@@
Scideator(smith2025scideator**) Sequential Refinement Human + Tool Hypothesis-Level Researchers iteratively refine idea facets with human guidance.
VirSci(johnson2024virsci) Sequential Refinement Peer agents Hypothesis-Level Iterative agent roles for generating, critiquing, voting ideas.
@@
AgentLab(schmidgall2025agentlaboratoryusingllm) Sequential Refinement Peer agents Hypothesis-Level Role-based agent pipeline from data to paper drafting.
@@
Digital Twin Parameterization(yang2025leveraging) Sequential Refinement Simulator + Peer agents Hypothesis-Level Multi-agent tuning of simulation parameters.
@@
IRIS(feng-etal-2025-iris) Branching Exploration Peer agents + Human Hypothesis-Level Uses MCTS for branching and incorporates both agent feedback and human steering.
@@
MC-NEST(liu2024mcnest) Branching Exploration Internal confidence Hypothesis-Level Explores reasoning paths via MCTS and evaluates them using internal metrics.
@@
MAGIC(wang2025magic) Branching Exploration Internal confidence Hypothesis-Level Uses multi-armed bandit scores to guide hypothesis branching.
@@
Monte Carlo Thought Search(kim2023montecarlothought) Branching Exploration Internal confidence Hypothesis-Level Applies combinatorial MCTS for catalyst design.
@@
Darwin Gödel Machine(zhang2025darwin) Branching Exploration Internal confidence Program-Level Uses open-ended evolution where agents rewrite and test their own code, preserving beneficial modifications.
Multi-Agent Debate(liang2023MAD) Branching Exploration Peer agents Hypothesis-Level Adversarial agent debates with referee judgment.
MOOSE-Chem2(yang2025moosechem) Parallel Refinemen Internal confidence Hypothesis-Level Samples many chemical hypotheses in parallel and scores them.
ChemCrow(bran2023chemcrow) Programmatic Planning External tool Program-Level Outputs synthetic routes executed via cheminformatics APIs.
CodeScientist(codescientist2024**) Programmatic Planning Simulator + Lab Execution Program-Level Writes lab scripts executed in simulator or robotics environments.
LLM-SR(shojaee2024llmsr) Programmatic Planning External tool Program-Level Generates symbolic equations that are evaluated and refined.
@@
Casevo(jiang2024casevocognitiveagentssocial) Sequential + Branching Simulator + Peer agents Hypothesis-Level Agent-based social simulation with opinion adaptation.
4.3.4 Takeaways and Limitations
Inference-time scaling techniques expand the effective search space of a single LLM, giving it more opportunities to explore, refine, and evaluate potential ideas. Table 1 provides a comparison of scientific inference-time scaling methods. However, this search is ultimately confined to one model’s internal distribution. For highly interdisciplinary or radically novel ideas, relying solely on a single model’s representation may be insufficient.

### v2

@@ 4.2.2 Peer LLMs as Verifiers
A growing line of work leverages LLMs as external evaluators, distinct from the generator itself. For example, FlexiVe (Zhong et al., 2025) implements a solve–detect–verify pipeline, while process-reward frameworks (Choudhury, 2025) and diverse verifier tree search use verifier models to score intermediate hypotheses and control expansions. Stability can be improved by ensemble verification (Wang et al., 2025a).
The CRPO (Ismayilzada et al., 2025a) exemplifies this approach by using an LLM’s general linguistic knowledge as a feedback signal for guiding logical consistency and language fluency. Similarly, the LLM-as-a-Judge (Zheng et al., 2023) paradigm formalizes this idea: one model (or ensemble (Chen et al., 2025a)) evaluates hypotheses produced by another, effectively operationalizing externalized language-model judgment as search heuristics.
More advanced methods first train specialized evaluators to capture problem-specific criteria before deploying them as verifiers. For example, CycleResearcher (Weng et al., 2025a) relies on a trained verifier model (CycleReviewer) to judge intermediate steps and direct search trajectories. Li et al. (2024c) similarly build a reward dataset and trains three reward models to guide idea generation dynamically, allowing fine-grained control over different evaluation dimensions. These approaches are often positioned as scalable, automated substitutes for human feedback, especially in domains where expert evaluation is costly. Peer evaluators are more flexible than self-confidence, since they may be domain-specialized and provide rationales.
However, they remain vulnerable to reward hacking: generators can learn to exploit systematic weaknesses in verifier heuristics, producing outputs that appear highly rated without being genuinely correct or useful. Recent studies in preference optimization and RLHF have shown that models can overfit to their evaluators, generating verbose, sycophantic, or superficially persuasive responses that maximize reward signals rather than substantive quality (Laidlaw et al., 2025; Miao et al., 2024). These findings highlight that while peer evaluators extend the evaluative horizon, their reliability depends on careful design to avoid exploitation and collapse of evaluation diversity. While effective as evaluators and filters, feedback sources differ fundamentally from collaborative multi-agent setups. As outlined in Section 5, multi-agent frameworks engage additional LLMs as co-creative partners, enabling them not only to assess but also to propose, contest, and refine ideas in a shared search process.
@@ 4.2.3 Simulators and Real Environments
Simulators serve as semi-grounded environments that test the outputs of LLMs in dynamic, feedback-rich contexts. These can be physical simulations, digital twins, or even agent-based societies. Simulators and real environments provide feedback through interactive trial-and-error, aligning closely with notion of online search(Russell and Norvig, 2020). Unlike offline planning, where a complete solution is computed in advance, online search emphasizes interleaving generation with execution: the agent proposes an action, observes the environment’s response, and adapts its trajectory accordingly. This makes simulators and lab-in-the-loop systems natural analogues, as they allow LLM-driven agents to iteratively refine hypotheses or strategies based on grounded feedback from physical laws, empirical trials, or synthetic environments.
Robin (Ghareeb et al., 2025) uses a “lab-in-the-loop” design, where hypotheses are tested via in silico simulators (or even robotic wet-labs), and the resulting outputs are looped back into the ideation process as reward signals. For example, a gene-editing strategy might be generated, simulated for efficacy, and then refined based on empirical response.
Similarly, AtomAgents (Ghafarollahi and Buehler, 2024) combines LLMs with physics-based alloy simulators: agents propose material compositions, query simulators for predicted properties like tensile strength, and prioritize candidates for further iteration based on performance metrics.
In more abstract contexts, CASEVO (Jiang et al., 2024) simulates agent societies debating political or scientific stances, tracking how persuasive interactions shift beliefs—modeling ideation in socially complex domains.
Recent robotics work demonstrates the coupling of simulators and real execution. SayCan (Ahn et al., 2022) grounds LLM-generated high-level commands in an affordance model and low-level robot controller, combining planning tools with real robot feedback. RoboTool (Xu et al., 2023b) extends this to creative tool use, where LLMs propose robot code tested in both simulators and physical hardware. Voyager (Wang et al., 2023a) highlights how open-ended exploration in Minecraft enables the synthesis of reusable tools, iteratively evaluated in a simulator to accumulate a growing library of skills.
On the evaluation side, Torne et al. (2024) analyze the fidelity of simulator-to-reality transfer, explicitly comparing robot manipulation policies in simulation versus physical execution. Similarly, Large-Scale Simulation of LLM-Driven Generative Agents stress-tests agentic LLMs in synthetic societies, showing how simulation feedback can approximate large-scale environment signals for long-running scientific ideation (Piao et al., 2025). As shown in the Scenario 7, the simulator evaluates stability and feasibility of candidate designs.
@@ 4.2.4 External Tools and APIs
External tools expand the feedback horizon by integrating APIs, online search, and domain-specific databases. The Toolformer architecture (Schick et al., 2023) equips LLMs with the capability to query external resources dynamically, enabling verification and augmentation of generated content. Chemical synthesis platforms like ChemCrow (Bran et al., 2023) leverage cheminformatics tools and external data sources to validate reaction pathways proposed by the model. Similarly, HuggingGPT (Shen et al., 2023) coordinates a wide range of specialized AI models hosted on the Hugging Face platform, allowing an LLM to act as a controller that routes tasks such as vision, speech, and text analysis to appropriate external tools, demonstrating how tool integration can expand an LLM’s capabilities beyond text-only reasoning.
Recent evaluations focus directly on LLM–tool interaction. MINT (Wang et al., 2024d) systematically studies multi-turn tool usage, showing how feedback from tool calls and natural language guidance improves decision-making. Voyager’s open-ended tool synthesis also bridges this category, where tools (skills, functions) are generated internally but then tested through simulators and APIs.
By incorporating external tools, LLMs move beyond isolated language generation toward grounded, multi-modal reasoning informed by up-to-date, external knowledge. This shift transforms them from pure text predictors into decision-theoretic agents augmented with specialized oracles. This feedback source and its alternatives are examined in Scenario 7.
@@ 4.2.5 Scientific Priors and Domain Rules
SCaSML (Fan et al., 2025) integrates physical laws such as energy conservation to constrain generation, ensuring outputs remain within scientifically plausible bounds without requiring explicit simulation feedback. Similarly, methods like R³V (Cheng et al., 2024) embed domain logic checkpoints to detect and correct illogical reasoning mid-generation. The Scenario 7 illustrates how scientific priors enforce fundamental constraints and ensure theoretical validity. Domain rules are most effective in well-formalized fields (e.g., physics, chemistry, mathematics) where symbolic constraints or laws are clear and non-negotiable. Their strength is reliability: they are resistant to reward hacking and scale well. However, their rigidity limits flexibility in ill-defined or emergent domains, where strict rules may exclude innovative or unconventional ideas.
@@ 4.2.6 Human-in-the-Loop Evaluation
Platforms like (He et al., 2024) integrate human expertise directly into the scientific discovery process by allowing researchers to iteratively guide and critique model-generated hypotheses. IdeaSynth (Pu et al., 2025) and Scideator (Radensky et al., 2024) provide interactive environments for hypothesis refinement, allowing researchers to blend and adapt elements from prior work. Chain-of-Table (Wang et al., 2024e) offers a mixed-initiative framework where human evaluation is combined with structured scientific tables and knowledge bases, balancing creativity with formal rigor. IRIS (Feng et al., 2025), though originally designed for self-reflection, also incorporates human evaluators in tasks requiring subjective judgment of novelty and coherence. These human-in-the-loop interactions enrich the feedback loop, ensuring that the discovery process not only leverages computational power for exploration but also incorporates experiential insights that are difficult to encode explicitly, ultimately leading to more reliable and interpretable scientific advancements.
Figure 7: Contrasting evaluation feedback sources for LLM-driven scientific ideation.
@@ 4.3 Levels of Abstraction
Inference-time scaling methods differ not only in their strategies but also in the level of abstraction at which search is conducted. From a theoretical perspective, abstraction provides a means of reducing or restructuring the problem space, thereby enabling broader exploration but complicating verification (Zucker, 2003; Nakar et al., 2024). Recent LLM-based methods explicitly operationalize this trade-off. For instance, Wang et al. (2024b) show that separating hypothesis generation at an abstract level from concrete program instantiation improves performance on reasoning tasks, while Singhal and Shroff (2024) demonstrate that concept-level scoring functions allow more efficient traversal of search spaces than surface-level metrics. Together, these studies suggest that higher levels of abstraction broaden the search space and bring the process closer to exploratory creativity, but at the cost of increased difficulty in assessing feasibility and correctness.
@@ 4.3.1 Hypothesis-Level Ideation
For instance, MAGIC (Xu et al., 2023a) generates speculative hypotheses in genomics and systems biology, ranking them by plausibility and novelty heuristics. VirSci (Johnson and Davis, 2024) applies multi-agent debate to refine hypotheses through peer critique. IRIS (Feng et al., 2025) adds internal reflection, restructuring ideas mid-generation to avoid incoherence or excessive derivativeness. These systems act as cognitive companions: their outputs are inspirations requiring downstream verification by humans or automated evaluators. An example of hypothesis-level ideation is illustrated in Scenario 8.
Figure 8: Illustration of hypothesis-level ideation in mathematical reasoning.
@@ 4.3.2 Program-Level Ideation
Program-Aided Language models (PAL) (Gao et al., 2023) exemplify this principle by translating natural-language reasoning into Python programs executed externally, outperforming direct chain-of-thought reasoning.HuggingGPT (Shen et al., 2023), Toolformer (Schick et al., 2023), and Explanations as Programs (Zhang and Huang, 2023) similarly teach models to output modular, debuggable tool pipelines. In science, LLM-SR (Shojaee et al., 2024) refines symbolic equations by querying solvers for constraint violations, ChemCrow (Bran et al., 2023) plans synthesis routes via cheminformatics APIs, and CodeScientist (Huang and Wu, 2024) generates executable lab protocols for simulators or robotic wet-labs.
Ideation at the program level enables the use of simulators and other external tools by integrating execution directly into the generative loop. An example of program-level ideation is illustrated in Scenario 9. This integration offers key benefits: programs are structured and interpretable, their execution provides robustness against unsupported claims, and hallucination is reduced as hypotheses must be empirically validated through execution.
Figure 9: Illustration of program-level ideation in mathematical reasoning.
@@ 4.3.3 Meta-Level Search and Self-Adaptation
A third, emerging frontier is meta-level search, where the object of optimization is not the hypothesis or program but the agent’s own configuration and interaction with its environment. Here, systems explore changes to their prompting strategies, exploration parameters, reward weightings, or even architectural components, adapting the search process itself. An example of meta-level ideation is illustrated in Scenario 10.
The Darwin Gödel Machine (Zhang et al., 2025a) exemplifies self-modification: instead of emitting domain-level programs, it evolves its own reasoning code to improve performance. More broadly, meta-search also includes adaptive inference strategies—e.g., adjusting beam widths, modifying exploration–exploitation trade-offs, or re-balancing between simulator calls and heuristic priors. Such mechanisms parallel meta-reasoning and utility-based agent adaptation, where the agent learns not only how to act, but how to search. Another notable example is the Meta-review agent in Google’s AI co-scientist system (Gottweis et al., 2025). This agent synthesizes feedback from various agents, identifies recurring issues, and generates comprehensive critiques and research overviews. The Meta-review agent compiles feedback across iterations to adjust the agents’ prompts, ensuring continual improvement.
@@
Figure 10: Illustration of meta-level ideation in mathematical reasoning.
@@
Table 1: Comparison of representative scientific inference time-scaling methods.
@@
Method Search Mechanism Feedback Source Abstraction Level Main Idea / Explanation
@@
Self-Refine (Madaan et al., 2023) Sequential refinement Internal confidence Hypothesis-level Iterative self-critique loop refining outputs.
@@
PANEL (Li et al., 2025d) Sequential refinement Internal confidence Hypothesis-level Stepwise reasoning refinement via critique generation.
@@
ResearchAgent (Baek et al., 2024) Sequential refinement Peer agents Hypothesis-level Simulates multi-agent peer review for idea improvement.
@@
CriticAL (Li et al., 2024b) Sequential refinement Internal confidence Hypothesis-level Applies statistical testing to critique and enhance reasoning.
@@
SCaSML (Fan et al., 2025) Sequential refinement Scientific rules Program-level Incorporates physical laws; refines via Monte Carlo correction.
@@
Robin (Ghareeb et al., 2025) Sequential refinement Simulator Hypothesis-level Refines ideas using simulated or in-lab feedback.
@@
IdeaSynth (Pu et al., 2025) Sequential refinement Human + Tool Hypothesis-level Human-guided iterative idea refinement.
@@
Scideator (Smith and Huang, 2025) Sequential refinement Human + Tool Hypothesis-level Joint human–tool refinement of idea components.
@@
AgentLab (Schmidgall et al., 2025) Sequential refinement Peer agents Hypothesis-level Role-based pipeline from data analysis to paper drafting.
@@
Digital Twin (Yang et al., 2025a) Sequential refinement Simulator + Peer agents Hypothesis-level Multi-agent tuning of simulation parameters.
@@
IRIS (Feng et al., 2025) Branching exploration Peer agents + Human Hypothesis-level MCTS-based branching with both agent and human feedback.
@@
MC-NEST (Rabby et al., 2024) Branching exploration Internal confidence Hypothesis-level Explores reasoning paths via MCTS with internal evaluation.
@@
MAGIC (Xu et al., 2023a) Branching exploration Internal confidence Hypothesis-level Guides hypothesis branching via multi-armed bandit scores.
@@
MCTS  (Sprueill et al., 2023) Branching exploration Internal confidence Hypothesis-level Applies combinatorial MCTS for catalyst design.
@@
DGM (Zhang et al., 2025a) Branching exploration Internal confidence Program-level Open-ended evolution: agents rewrite and test their own code.
@@
Casevo (Jiang et al., 2024) Sequential + Branching Simulator + Peer agents Hypothesis-level Agent-based social simulation with adaptive opinion modeling.
MOOSEChem2 (Yang et al., 2025b) Parallel refinement Internal confidence Hypothesis-level Parallel sampling of chemical hypotheses with internal scoring.
ChemCrow (Bran et al., 2023) Programmatic planning External tool Program-level Designs synthesis routes executed via cheminformatics APIs.
CodeScientist (Huang and Wu, 2024) Programmatic planning Simulator + In-lab run Program-level Writes lab scripts run in simulators or robotic labs.
LLM-SR (Shojaee et al., 2024) Programmatic planning External tool Program-level Generates and refines symbolic equations through tool feedback.
4.4 Takeaways and Limitations
Inference-time scaling techniques expand the effective search space of a single LLM, giving it more opportunities to explore, refine, and evaluate potential ideas. Table 1 provides a comparison of scientific inference-time scaling methods. Empirical evidence from Monte Carlo Thought Search(Sprueill et al., 2023) demonstrates the impact of test-time scaling and structured search on scientific ideation quality. Using a domain-specific reward function, the study shows that increasing search depth and breadth substantially improves outcomes compared to single-pass prompting. While standard CoT achieves low rewards (2.04–2.27), adding self-consistency yields moderate gains (4.04–6.38). In contrast, structured search methods such as Tree-of-Thought (breadth-first search) achieve much higher rewards (9.91–13.8) by evaluating hundreds of candidate prompts (NP = 253, depth = 5). Monte Carlo Thought Search further improves performance (12.47–15.6) by deeper and more adaptive exploration (NP $\approx$ 301, depth $\approx$ 9). These results indicate that test-time scaling enables exploratory creativity by systematically traversing the ideation space, though gains come at the cost of increased computational expense and reliance on evaluator quality.